OpenStudy (anonymous):

7/9 divided by 21/27

OpenStudy (anonymous):

that is 1

OpenStudy (maheshmeghwal9):

\[\color{red}{\frac{7}{9}\div \frac{21}{27}.}\]\[\color{blue}{\frac{7}{9}\times \frac{27}{21}.}\]\[\color{green}{\frac{3}{3}=1.}\]

OpenStudy (maheshmeghwal9):